OpenStudy (anonymous):

Six boys or four men can complete a piece of work in 24days. In how many days will 3 boys and 10 men together complete the same piece of work?

OpenStudy (amistre64):

1 boy = 1/6 of the job 1 man = 1/4 of the job 3(1/6) + 10(1/4) = K 1/K = time to complete

OpenStudy (amistre64):

something along those lines

hero (hero):

"six boys OR four men"... is that correctly expressed?

OpenStudy (anonymous):

s

OpenStudy (amistre64):

3(1/6) + 10(1/4) = K 1/2 + 5/2 = K 3 = K 24(1/3) = 8

hero (hero):

@amistre64 do you agree with the wording of this problem? Using "or"?

OpenStudy (anonymous):

it better be OR if it is AND you cannot do it

OpenStudy (anonymous):

suppose you say Six boys AND four men can complete a piece of work in 24days for all you know the boys do all the work and the four men go out and drink beer all day

hero (hero):

I've never seen a working together problem with "or" in it.

OpenStudy (amistre64):

i see no issue with the use of "or" ; we are still able to define the rate of a boy or a man with the information given: 4m = 1 job = 24 days m = 1/4 job = 6 days worth of work 6b = 1 job = 24 days b = 1/6 job = 4 days worth of work

OpenStudy (amistre64):

10(6) + 3(4) = 72 days of work packed into the job; which is 3 jobs worth

OpenStudy (amistre64):

they are working 3 times as fast, therefore a 24 day job is done in 1/3 the time

OpenStudy (anonymous):

i got it.see 10/96+3/144=8

OpenStudy (amistre64):

thatll work too :)

OpenStudy (anonymous):

then wat ur trying to say

OpenStudy (amistre64):

i said what i was trying to say ... i was reasoning thru the process

OpenStudy (anonymous):

means

OpenStudy (anonymous):

grt @amistre64

OpenStudy (anonymous):

wat r u doing now

hero (hero):

@amistre64 @satellite73

OpenStudy (amistre64):

thats a fine method too :)

hero (hero):

It's @satellite73 method. He still hasn't told me whether or not he's the originator of that setup.

OpenStudy (amistre64):

i heard he got it from some guy in Sparta ...

OpenStudy (amistre64):

but i just cant trust the voices in my head lol

hero (hero):

I see you still got jokes.

OpenStudy (anonymous):

grt all who replied my question

OpenStudy (amistre64):

good luck anala :)

OpenStudy (anonymous):

not only did i invent it, but i copyrighted it as well, and get a royalty every time it is used my invoice is in the mail

hero (hero):

Awesome :) Share the wealth please.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

hey how m=2.4

hero (hero):

Good question.

hero (hero):

b and m represent something other than number of boys and number of men. I admit I didn't define it well.

hero (hero):

3 is the number of boys and 10 is the number of men.

hero (hero):

Actually @amistre64

hero (hero):

Let me post another solution I think is correct.

hero (hero):

b = 4 m = 6 \[\frac{(3b)(10m)}{3b + 10m} = x\] \[\frac{(12)(60)}{12 + 60} = 10\]

hero (hero):

b = days each boy worked m = days each man worked

hero (hero):

The solution I posted earlier...there was a fundamental flaw in it. I computed the wrong number the first time.

hero (hero):

If it's wrong, please at least explain why it is wrong.

hero (hero):

I think I'm beginning to see why it is wrong.

hero (hero):

There's no way 3 boys can complete the work in 12 days. That would be faster than it takes the six to complete in 24 days.

hero (hero):

:(

hero (hero):

Maybe I'm looking at it the wrong way. Maybe the 3 boys could only do 12 total days worth of work because there are only three boys.

OpenStudy (amistre64):

it takes 24 days for 6 boys to complete 1 job; it takes 1 day for 6 boys to complete 1/24 of the job; each boy does 1/6 of 1/24 of a job in 1 day; b = 1/144 of the job per day it takes 24 days for 4 men to complete 1 job; it takes 1 day for 4 mens to complete 1/24 of the job; each man does 1/4 of 1/24 of a job in 1 day; m = 1/96 of the job per day

OpenStudy (amistre64):

3(1/144) + 10(1/96) is the amount of work done in 1 day

OpenStudy (amistre64):

1/8 of the work is done in 1 day; therefore it takes 8 days to do one job

hero (hero):

Well my alternative method attempts have failed.

hero (hero):

i think inverse proportions are involved with this. That's why you used k as the variable.

OpenStudy (amistre64):

you are not calculating the appropriate b,m,workday ratio is all

hero (hero):

I solved enough working together problems to know that there's alternative ways to solve them.

OpenStudy (amistre64):

there are :)

hero (hero):

I'll figure it out. Thanks for the extra tips.

OpenStudy (amistre64):

youre welcome